# Radial Fast Entangling Gates Under Micromotion in Trapped-Ion Quantum Computers

### v1

@@ Abstract
Micromotion in radio-frequency ion traps is generally considered detrimental for quantum logic gates, and is typically minimized in state-of-the-art experiments. However, as a deterministic effect, it can be incorporated into quantum control frameworks aimed at designing high-fidelity quantum logic controls. In this work, we demonstrate that micromotion can be beneficial to the design of fast gates utilizing the radial modes of a two-ion crystal, particularly in the sub-trap-period regime where high-fidelity control sequences are identified with operation times ranging from hundreds of nanoseconds to microseconds. Through analysis of select fast gate solutions, we uncover the physical origin of micromotion enhancement and further study the induced gate error under experimental noises and control imperfections. This analysis establishes the feasibility of realising high-fidelity entangling gates in hundreds of nanoseconds using the micromotion-sensitive radial modes of trapped-ion crystals.
@@ I Introduction
Trapped ions are a leading platform for quantum computing [Cirac1995a, Bruzewicz2019e, Drmota2024], quantum networking [Stephenson2020, Drmota2023a, OReilly2024], quantum metrology [Marshall2025, Nichol2022], and tests of fundamental physics [Cairncross2017, Carney2021]. For quantum information processing, trapped ion devices are advantageous due to their long coherence times [Langer2005, Harty2014, Wang2021a], demonstrated high-fidelity operations [Harty2014, Ballance2016, Srinivas2021a, Ryan-Anderson2024, Moses2023d, Leu2023, Cai2023, Weber2024], and high qubit connectivity mediated by crystal ion vibrations [Wright2019, Debnath2016, Figgatt2019, Hou2024, Grzesiak2020, Chen2024] or physical shuttling of ion qubits [Kielpinski2002, Wan2019, Kaushal2020b, Pino2021, Moses2023d, Mordini2025, Akhtar2023]. State-of-the-art ion-trapped systems rely on radio-frequency (RF) Paul traps [Paul1990a], where a strong AC driving RF signal leads to a deterministic modulation of the ion-crystal’s harmonic motion, known as intrinsic RF micromotion [Leibfried2003b]. There is a growing body of work suggesting the importance of including micromotion in quantum logic gate design [Shen2014, Wang2015, Bermudez2017, Wu2021, Ratcliffe2020], including recent work on entangling gate schemes that leverage intrinsic micromotion to improve entangling gate operations [Bermudez2017, Ratcliffe2020]. This work contributes to the latter by demonstrating that intrinsic micromotion offers advantages in the design of fast and high-fidelity entangling gates between trapped-ion qubits.
RF micromotion is not temporally resolved in typical quantum logic gates in trapped-ion qubits, where the driving fields induce Rabi oscillations on timescales much longer than the RF period. Fast single-qubit gates have been demonstrated with durations of tens of picoseconds [Madsen2006, Campbell2010b, Mizrahi2013c, Mizrahi2014] and machine-designed fast two-qubit gates have been demonstrated with durations of microseconds [Schafer2018b, Wong-Campos2017a]. Fast gate schemes using stepwise modulation of the driving laser have demonstrated the highest fidelity, but are limited by photon scattering and the breakdown of the Lamb-Dicke regime [Schafer2018b]. Fast gate schemes using sequences of picosecond pulses to generate state-dependent momentum kicks (SDKs) are limited by the fidelity of those SDKs, which are more experimentally challenging than MHz modulation [Wong-Campos2017a]. However, fast gates based on SDKs are not limited by spontanteous emission or to the Lamb-Dicke regime [Garcia-Ripoll2003c, Garcia-Ripoll2005c, Zhu2006b, Bentley2013, Taylor2017b, Gale2020, Savill-Brown2025b]. Proposals have shown that they scale to larger ion crystals with ‘all-to-all’-connected coupling [Mehdi2021e, Savill-Brown2025, Savill-Brown2025b], and can be used in flexible environments such as mixed species ion crystals [Mehdi2025] and microtrap arrays [Ratcliffe2018, Mehdi2020b, Wu2020b].
Micromotion is typically neglected in proposed fast gate protocols, i.e. by assuming excitation along the RF null of a linear Paul trap. However, individual ion addressing of ‘micromotion-free’ axial modes is challenging due to laser alignment errors and mechanical instabilities. Some experimental configurations require two-photon driving from carefully oriented Raman beams [Mehdi2025, Savill-Brown2025b] which makes coupling to radial modes harder to avoid. These issues are compounded when considering larger ion arrays and scalable geometries. Gates that operate using the micromotion-sensitive radial modes of ion crystals are, therefore, highly desirable. In this work, we find that in some parameter regimes, the high-frequency modulation can actually be used to improve gate speed and fidelity.
@@
Previous work on the development of fast gates has either considered micromotion-free systems or assumed that the excitation is synchronized to the RF drive. A study by Ratcliffe et al. [Ratcliffe2020] demonstrated that micromotion can improve the speed of two-qubit fast gates. This enhancement is achieved by locking laser pulses to a specific phase of the RF drive, with a larger micromotion modulation leading to a greater potential benefit. While all fast gate implementations will require precise control over the timing of pulses relative to the RF drive, this phase-locking also places the strong constraint that the RF frequency must be a multiple of the laser repetition rate. Furthermore, this restriction ignores all gate solutions that leverage the full temporal structure of the RF drive, which opens a many new degrees of freedom.
@@
This work extends the findings of Ratcliffe et al. [Ratcliffe2020] in a two-ion chain. We lift the constraint that locks the phase of the pulses to the RF drive, which enables more experimentally feasible parameter regimes. We further expand previous results by employing a more flexible optimisation method, the Generalised Pulse Group (GPG) scheme [Gale2020], which we extend to incorporate the full temporal structure of the RF drive. Our results demonstrate that high-fidelity entangling gate solutions can be found with operation times faster than a radial trapping period, across a wide range of micromotion environments. We demonstrate that regimes exist where larger amplitude micromotion improves both gate speed and fidelity, which we analyse in terms of the deviation from the secular (micromotion-free) limit. We further investigate the experimental feasibility and susceptibility of micromotion-enabled fast gates to various error sources. We find that fidelities exceeding $0.9$ may be achieved using previously demonstrated SDK fidelities, and that gate errors approaching $10^{-5}$ are possible with high-fidelity SDKs. Our results establish the viability of achieving MHz-rate multi-qubit quantum logic utilizing the radial modes of a linear Paul trap configuration.
Figure 1: Impulsive fast two-qubit gates using the transverse (radial) modes of a two-ion crystal. (a) i. Ions are subjected to SDKs using focussed optical beams (shown in red) transverse to the null of the RF trap (dashed line). ii. Addressing both ions with SDKs excites the in-phase (common-motional) mode for same-spin two-qubit states, and the out-of-phase (stretch) mode for different-spin two-qubit states. (b) First-order stability region for the trapping parameters $(a,q)$ along the $x$-axis (blue shaded region) and the $y$-axis (red shaded region). The overlap between the two regions corresponds to stable trapping in all dimensions (hatched region). Trapping parameter values along the $x$-axis chosen to maintain the same trapping potential in the secular limit indicated with dots (diamonds) for an RF drive of $40\omega_{0}$ ($20\omega_{0}$), where $\omega_{0}$ is the secular frequency of the RF trap. (c) i. Example dynamics of fast gates of duration $\approx 0.7(2\pi/\omega_{0})$ for regimes of low-amplitude micromotion ($q_{x}=0.01$) and high-amplitude micromotion ($q_{x}=0.5$). ii. Arrival and magnitude of SDKs within the RF cycle in the high-micromotion regime, $q_{x}=0.5$, illustrating state-dependent impulses occur at various points in the RF cycle. iii. Histogram of the RF phases at the time of each SDK ($t_{k}$), i.e. $\phi^{(k)}=\phi_{\rm RF}+\Omega_{\rm RF}t_{k}$.
II Background Theory: Quantum dynamics in Radio-frequency traps
Radio frequency (RF) ion traps are generated by a combination of DC and AC electric fields, which create an effective trapping potential with static (DC) and dynamic (AC) contributions. In this work we focus on the linear Paul trap configuration: static trapping along the linear axis of the trap ($z$) and dynamic trapping in the xyx-y plane. We further assume excitation of the ion crystal purely along the $x$ axis, such that we can consider a one-dimensional model for the potential energy of a single trapped ion:

### v2

@@ Abstract
Micromotion in radio-frequency ion traps is generally considered detrimental for quantum logic gates, and is typically minimized in state-of-the-art experiments. However, as a deterministic effect, it can be incorporated into quantum control frameworks aimed at designing high-fidelity quantum logic controls. In this work, we demonstrate that micromotion can be beneficial to the design of fast entangling gates utilizing the radial modes of a two-ion crystal, particularly in the sub-trap-period regime where high-fidelity control sequences are identified with operation times ranging from hundreds of nanoseconds to microseconds. Through analysis of select fast gate solutions, we uncover the physical origin of micromotion enhancement and further study the induced gate error under experimental noises and control imperfections. This analysis establishes the feasibility of realising high-fidelity entangling gates in hundreds of nanoseconds using the micromotion-sensitive radial modes of trapped-ion crystals.
@@ I Introduction
Trapped atomic ions have a number of advantages for quantum information processing [13, 8, 17]: exceptional coherence times of qubits encoded in electronic states [34, 25, 71], high-fidelity quantum logic operations [25, 2, 63, 55, 47, 36, 10, 72], and strong qubit connectivity [14, 74, 28, 32, 51, 47]. State-of-the-art trapped-ion processors rely on radio-frequency (RF) Paul traps [50], where a strong AC driving field induces a deterministic modulation of the ion-crystal’s harmonic motion, known as intrinsic micromotion [35]. Micromotion is typically neglected in the design of conventional quantum logic gates, such as the Mølmer-Sørenson entangling mechanism [61], which operate on adiabatic timescales with respect to the trapping fields [8]. However, the effects of micromotion become increasingly important as gate speeds are pushed towards the non-adiabatic regime, where the control fields are adjusted at timescales comparable to or faster than the RF drive [53].
Non-adiabatic ‘fast’ entangling gates use high-power laser pulses to implement two-qubit operations on timescales comparable to the trapping period [22, 64]. This can be achieved either through amplitude modulation of the driving laser within the Lamb-Dicke regime [58], or through sequences of state-dependent momentum kicks (SDKs) delivered by ultrafast $\pi$-pulses [21]. The latter is limited by the fidelity of those SDKs [73] which are more experimentally challenging than MHz modulation. However, SDK-based fast gates with machine-designed pulse sequences [20] have a number of advantages: they support temperatures beyond the Lamb-Dicke regime [67], scale to larger ion crystals with all-to-all connectivity [41, 57, 56], operate between ions in neighbouring traps [54, 40, 76], and support mixed-species ion pairings [42].
@@
The purpose of this work is to explore how micromotion can be exploited in the design of high-fidelity SDK-based fast gates utilizing the radial modes of a trapped-ion crystal. A study by Ratcliffe et al. [53] demonstrated the existence of regimes where micromotion can improve the speed of two-qubit fast gates by locking laser pulses to a specific phase of the RF drive, with larger micromotion amplitudes leading to greater potential benefit. However, this phase-locking also places the strong constraint that the RF frequency must be a multiple of the laser repetition rate. Furthermore, this restriction ignores all gate solutions that leverage the full temporal structure of the RF drive, which opens up many new control degrees of freedom.
@@
This work extends the findings of Ratcliffe et al. [53] to a phase-unlocked regime, where SDKs are no longer constrained to arrive at a fixed point in the RF cycle. Using numerical optimization over high-dimensional pulse sequences [20], we identify high-fidelity entangling gate solutions with operation times comparable to and faster than a single radial trapping period across a wide range of micromotion environments. We find that larger amplitude micromotion can improve both gate speed and fidelity; the latter by up to two orders of magnitude. We analyse the physical origin of this enhancement in terms of the deviation from the secular (micromotion-free) limit, and study the susceptibility of these gates to experimental error sources, finding that fidelities exceeding $0.9$ may be achieved using previously demonstrated SDK fidelities, and that gate errors approaching $10^{-5}$ are possible with high-fidelity SDKs [37]. Our results establish the viability of achieving MHz-rate entangling gates in environments with significant micromotion, such as the radial modes of a Paul trap.
II Theory of fast gates under micromotion
In this section we outline the SDK-based fast gate formalism in the presence of micromotion. This extends the original fast gate formalism developed in Ref. [21] to incorporate the full time-dependent dynamics of the RF trap (see Appendix A). The key result of this section is a set of condition equations for high-fidelity entanglement [Equations (8a)-(8c)], which highlight micromotion-induced corrections to the gate dynamics. These results are valid for arbitrary length chains, though we will focus on the $N=2$ case for the remainder of this work.
[…]
II.1 State-dependent kicks
Fast gates rely on the use of broadband (picosecond to nanosecond) $\pi$-pulses that impulsively transfer internal-state-dependent momentum to each of the addressed ions as illustrated in Fig. 1. Each of these SDKs can be realised by addressing each ion with a laser pulse of area $\pi$, resonant with the qubit frequency with effective wavevector $k$. A $\pi$-pulse incident on the $j$-th ion is described by the unitary:
where $\phi_{L}$ is the phase of the laser electric field and $\hat{x}^{(j)}(t)$ describes the deviation of the $j$-th ion’s position from the RF null (see Fig. 1(a)).Here we are working in an interaction picture with respect to the motional dynamics of the ion crystal in a time-dependent RF potential (see Appendix A):
$V_{\rm RF}(x,t)=\frac{m\Omega_{\rm RF}^{2}}{8}\left[a_{x}-2q_{x}\cos(\Omega_{\rm RF}t+\phi_{\rm RF})\right]x^{2}\,$, (2)
characterised by the dimensionless parameters $a_{x}$ and $q_{x}$.
In this work we consider a laser-phase-insensitive ‘paired’ SDK protocol, where each SDK is composed of a pair of counter-propagating $\pi$-pulses incident on each ion (e.g. by retroreflecting each pulse [5, 52]). If the phase of the two pulses in each pair are the same (which can be ensured by splitting them from a $2\pi$ parent pulse), the laser-phase dependence of the paired SDK vanishes:
$\hat{U}_{\rm{SDK}}(k,t)$ $=\prod_{j=A,B}\hat{U}_{-k,\phi}^{j}(t)\hat{U}_{k,\phi}^{j}(t)$ (3)
$=\exp\left\{2ik(\hat{x}^{A}(t)\hat{\sigma}_{z}^{A}+\hat{x}^{B}(t)\hat{\sigma}_{z}^{B})\right\}\,.$ (4)
The action of this unitary on the two-qubit system is illustrated in Fig. 1(b). We will assume a paired SDK scheme for the remainder of this work, though the extension to unpaired $\pi$-pulses is straightforward [73, 56].
II.2 The fast gate unitary
$\hat{U}_{\rm G}$ $=\mathcal{T}\prod_{j=1}^{\mathcal{N}}\hat{U}_{\rm{SDK}}(z_{j}k,t_{j})\,$, (5)
with $\mathcal{T}$ denoting the time-ordering operator. Here we have allowed for switching the direction of each SDK between +k+k and k-k by the introduction of the parameter $z_{j}=\pm 1$ for the $j$-th SDK. Applying the Baker-Campbell-Hausdorff lemma, $\hat{U}_{\rm G}$ may be expressed as the product of two commuting unitaries, i.e. $\hat{U}_{\rm G}=\hat{U}_{2}\hat{U}_{1}$, where the first unitary describes an effective spin-spin interaction between the addressed qubits:
$\hat{U}_{1}=\exp\left\{i\Theta\hat{\sigma}_{z}^{(A)}\otimes\hat{\sigma}_{z}^{(B)}\right\}\,$, (6)
$\hat{U}_{2}=\bigotimes_{\alpha}\hat{D}_{\alpha}\left\{\Delta\xi_{\alpha}\sum_{j=A,B}b_{\alpha}^{(j)}\hat{\sigma}_{z}^{(j)}\right\}\,$, (7)
where $\Delta\xi_{\alpha}$ is the amplitude of residual spin-dependent displacement of the $\alpha$-th motional mode at the end of the gate operation, and $b_{\alpha}^{(j)}$ is the normalized coupling of the $j$-th ion to the $\alpha$-th motional mode. For the two-ion system, $b_{\rm CM}^{(1)}=b_{\rm CM}^{(2)}=1/\sqrt{2}$ for the center-of-mass (CM) mode and $b_{\rm SR}^{(1)}=-b_{\rm SR}^{(2)}=1/\sqrt{2}$ for the stretch (SR) mode.
Provided that the motional states are decoupled from the internal states at the conclusion of the gate operation (i.e. $\Delta\xi_{\alpha}=0$), the gate unitary takes the form of a ZZ entangling operation $U_{\rm ZZ}(\theta)=e^{i\theta\hat{\sigma}_{z}^{(A)}\otimes\hat{\sigma}_{z}^{(B)}}$]. Maximum entanglement is achieved for $\Theta_{\rm targ}=\pm\pi/4$. The key challenge to designing a high-fidelity operation is finding a pulse sequence $\{z_{k},t_{k}\}$ that satisfies the maximally-entangling condition, $\Theta=\pm\pi/4$, and restores all motional modes $\Delta\xi_{\alpha}=0$]. We describe a machine-design approach to address this problem in the following section.
The timings and directions of the SDK sequence enter the gate unitary through the quantities $\Theta$ and $\Delta\xi_{\alpha}$, which respectively describe the relative phases between the eigenstates of $\hat{\sigma}_{z}^{(A)}\otimes\hat{\sigma}_{z}^{(B)}$ and the residual spin-dependent displacement amplitude of the $\alpha$-th motion mode. Expressing the residual displacement in terms of the dimensionless mode quadratures, i.e. $\Delta X_{\alpha}=[\Delta\xi_{\alpha}u_{\alpha}^{*}(t_{\rm g})+\Delta\xi_{\alpha}^{*}u_{\alpha}(t_{\rm g})]/\sqrt{2}$ and $\omega_{\alpha}\Delta Y_{\alpha}=[\Delta\xi_{\alpha}\dot{u}_{\alpha}^{*}(t_{\rm g})+\Delta\xi_{\alpha}^{*}\dot{u}_{\alpha}(t_{\rm g})]/\sqrt{2}$, the following expressions for these quantities are derived in Appendix B: